Bank will give you after 5-years $1000, 000 with considering 12%-interest. For
this situation evaluate the present, annual, & future values at simple & compound
interest rates.
1
Expert's answer
2020-05-24T18:09:52-0400

FV simple & compound:$1000, 000

PV  simple:

"FV=PV\times(1+i\times n)"

"1 000 000=PV\times(1+0.12\times 5)"

PV=625 000


PV compound:

"FV=PV\times(1+i)^n"

"1 000 000=PV\times(1+0.12)^5"

PV=567 426.86


annual

"FV=625 000\times(1+0.12\times1)=700 000"

"FV=567 426.86\times(1+0.12)^1=635518.08"